# LEO-VL: Towards 3D Vision-Language Generalists via Data Scaling with Efficient Representation

### v1

@@ 3.3 Post-training
Comprehensive instruction tuning alone may be insufficient to ensure model robustness, as evidenced by prior studies in 2D-VL models [43, 66, 87]. Given the pronounced risk of overfitting in 3D-VL models [22, 35], it is essential to design an effective post-training objective to enhance their robustness.
To this end, we propose SceneDPO, a novel learning objective for the post-training of 3D-VL models. First, we follow DPO to adopt a loss term that contrasts between positive answer and negative answer . Motivated by the issue of visual ignorance [35] in 3D-VL models, where answers are predicted by overfitting to text without utilizing scene information, we introduce a loss term that contrasts between positive scene and negative scene . This discourages model from predicting current answer when conditioned on irrelevant scenes. To mitigate degradation of positive answers, we also incorporate a negative log-likelihood loss term , which proves critical in prior works [50, 62, 87]. Given a training tuple , the overall loss is defined as follows:
@@ 4 Experiment
In this section, we first compare LEO-VL against state-of-the-art models on a variety of 3D QA benchmarks in Section 4.1, highlighting its advantages in both efficiency and accuracy. In Section 4.2, we present ablation studies on model design to show the effectiveness of CFG. In Section 4.3, we analyze various training data configurations to reveal the influence of task diversity, scene domain coverage, data quality, and data scale on 3D-VL generalist learning. In Section 4.4, we further explore the effect of post-training objective based on LEO-VL.
@@
We initialize the 2D visual encoder and LLM with Qwen2.5-VL-7B-Instruct [4]. The learnable parameters include position embedding, anchor embedding, and LoRA parameters [30] of the LLM, which amount to 66M in total. We set the voxel size to 0.2 meters and retain up to 750 scene tokens of CFG. We train LEO-VL on our comprehensive 3D-VL dataset for 5 epochs, which takes 2 days with 8 NVIDIA A100 80G GPUs. We adopt AdamW optimizer [51] with a base learning rate at , scheduled with linear warmup and cosine decay. We provide detailed hyperparameters in Appendix.
@@
We primarily evaluate LEO-VL on 3D QA benchmarks considering the limited availability of standardized benchmarks for captioning, planning, and dialogue. Existing object captioning benchmarks [14, 2] are excluded in the main paper (see results in Appendix C), as their captions are derived from object grounding texts with simplistic patterns that cannot reflect natural descriptive capabilities. Our evaluation includes ScanQA [3] for general QA in 3D scenes, SQA3D [56] for situated QA, and MSQA [46] for situated QA across multiple scene domains (i.e., ScanNet, 3RScan, and ARKitScenes). We also incorporate Beacon3D [35], which offers robust metrics for 3D QA evaluation. Evaluation metrics include n-gram metrics (CIDEr and BLEU-4) for ScanQA, exact-match accuracy (EM and EM-R) for SQA3D, and GPT-Score for MSQA and Beacon3D.
@@ 4.1 Comparison with State-of-the-art Models
Model | Scene (#tokens) | ScanQA (val) | SQA3D (test) | MSQA (test) | Beacon3D | MSQA (test) | MSQA (test) | |||
C | B-4 | EM | EM-R | GPT-Score | Case | Obj. | GPT-Score | GPT-Score | ||
ScanQA [3] | Query (256) | 64.9 | 10.1 | 47.2 | - | - | - | - | - | - |
3D-LLM [29] | Query (32) | 74.5 | 12.9 | 49.8 | - | - | - | - | - | - |
PQ3D [118] | Query (80) | - | - | 47.1 | - | - | 35.9 | 4.2 | - | - |
DSPNet [52] | Query (256) | - | - | 50.4 | - | - | - | - | - | - |
3D-LLaVA [21] | Query (100) | 92.6 | 17.1 | 54.5 | 56.6 | - | - | - | - | - |
3D-VisTA [117] | Object (80) | 69.6 | 10.4 | 48.5 | - | - | 35.3† | 8.1† | - | - |
LEO [34] | Object (60) | 101.4 | 13.2 | 50.0 | 52.4 | 56.7† | 45.2 | 7.5 | 50.4† | 59.7† |
SceneVerse [39] | Object (80) | - | - | 49.9 | - | - | 40.5 | 4.7 | - | - |
Chat-Scene [32] | Object (200) | 87.7 | 14.3 | 54.6 | 57.5 | - | 45.8† | 7.8† | - | - |
Inst3D-LMM [99] | Object (200) | 88.6 | 14.9 | - | - | - | - | - | - | - |
Scene-LLM [26] | Voxel (0.18m) | 80.0 | 11.7 | 53.6 | - | - | - | - | - | - |
LLaVA-3D [115] | Voxel (3096) | 91.7 | 14.5 | 55.6 | 57.6 | - | - | - | - | - |
Video-3D LLM [111] | Video (32 frames) | 100.5 | 16.3 | 57.7 | - | - | - | - | - | - |
GPT4Scene [71] | Video (32 frames) | 96.3 | 15.5 | 59.4 | 62.4 | - | - | - | - | - |
LEO-VL | Grid (750) | 97.6 | 14.8 | 59.7 | 62.6 | 58.2 | 59.5 | 18.2 | 51.4 | 64.2 |
@@
For comparison, we include state-of-the-art 3D-VL models across four categories: (1) query-based methods, ranging from early work like ScanQA [3] to recent 3D-LLaVA [21]; (2) object-centric methods, from 3D-VisTA [117] to Inst3D-LMM [99]; (3) voxel-based methods, such as Scene-LLM [26] and LLaVA-3D [115]; and (4) video-based methods, including Video-3D LLM [111] and GPT4Scene [71]. We also report each model’s scene representation and its associated cost. For Scene-LLM, only the voxel size (0.18 meters) is available, which we estimate to yield a similar number of tokens as LLaVA-3D (voxel size 0.2 meters). For video-based methods, we list the number of frames, noting that 32 frames typically amount to over 6k scene tokens.
@@ Results.
As shown in Section 4.1, LEO-VL achieves state-of-the-art performance on most 3D-VL benchmarks with a low representation cost. While video-based models show competitive performance, their representations involve thousands of tokens and limit data scalability, hindering the acquisition of generalist capabilities through large-scale training. In contrast, the efficient representation of LEO-VL enables 3D-VL data scaling for consistently strong performance across domains including ScanNet, 3RScan, and ARKitScenes. These results underscore the advantages of our efficient representation in perception capability and data scalability.
@@ 4.2 Model Analysis
In Fig. 4, we present statistics of voxel tokens and CFG tokens on ScanNet. In addition to the distribution of token counts, we report two metrics: compression rate, defined as the ratio of CFG tokens to raw voxel tokens before condensation; and preservation rate, defined as the proportion of voxels retained by the CFG tokens. The results show that CFG achieves an average compression rate of about and an average preservation rate close to 99%. This demonstrates the high efficiency of our CFG representation.
@@
We provide a joint visualization of efficiency and accuracy in Fig. 3. Accuracy is measured using the EM on SQA3D, the metric with the most available reference results. As shown in Fig. 3, LEO-VL achieves a Pareto optimum between efficiency and accuracy. To investigate the efficiency of CFG, we ablate the representation with two alternatives on the ScanNet subset: voxel-full, which retains all voxel tokens without condensation; and voxel-sample, which downsamples voxel tokens to match the token count of CFG. The results in Fig. 3 show that compared to voxel-full, voxel-sample suffers a drop in accuracy, while CFG even slightly improves the accuracy. These results suggest that high accuracy in 3D-VL tasks can be achieved with significantly lower token costs, as enabled by our efficient CFG representation.
@@
A key advantage of LEO-VL lies in the disentanglement of 3D spatial modeling across vertical and horizontal dimensions. To assess the effectiveness of position embeddings, we train LEO-VL on the ScanNet subset under two ablated settings: “w/o z-pos”, where the z-axis position embedding is removed; and “w/o xyz-pos”, where both z-axis and xy-plane position embeddings are removed. As shown in Fig. 5, both types of position embeddings play critical roles in 3D-VL tasks. Removing the z-axis position embedding leads to a general performance drop, while removing the xy-plane position embedding especially degrades performance on SQA3D. These results validate the effectiveness of our design in 3D spatial modeling.
@@ 4.3 Data Analysis
We conduct ablative studies on data to answer the following questions: (1) What is the effect of captioning, planning and dialogue tasks in 3D-VL generalist learning? (2) What is the benefit of scaling across diverse scene domains? (3) What happens when we naively scale up 3D-VL data without considering data quality? (4) How effective is data scaling for LEO-VL with curated data?
@@
Based on LEO-VL (ScanNet), we ablate the training data with different task configurations: “w/o cap” removes both object captioning and scene captioning tasks, while “w/o pl+dl” removes planning and dialogue tasks. As shown in Fig. 6, both task categories contribute meaningfully to the generalist capabilities of LEO-VL. Captioning proves particularly beneficial for SQA3D, whereas planning and dialogue tasks are more influential for ScanQA. We further perform a qualitative test by prompting “w/o cap” model with a scene captioning task. The model struggles to generate meaningful scene descriptions, likely due to the lack of training in captioning. These findings underscore the importance of task diversity for both acquiring task-specific skills and enhancing cross-task performance, thereby playing a critical role in achieving generalist capabilities.
@@
To assess the impact of scaling across scene domains, we introduce an additional checkpoint trained on ScanNet and 3RScan (Scan+3R), complementing existing LEO-VL and LEO-VL (ScanNet). As shown in Section 4.3, compared to “ScanNet”, “Scan+3R” yields significant improvements on 3RScan and ARKitScenes, with only a slight performance drop on ScanNet. Further, compared to “Scan+3R”, “All” shows substantial gains on ARKitScenes. These results suggest that the performance gains may primarily stem from exposure to corresponding domains, emphasizing the importance of a wide domain coverage for robust cross-domain generalist performances.
@@
Based on the results of “Scan+3R”, we explore the outcome of naively scaling up 3D-VL data without considering data quality. Specifically, we augment our default subsets on ScanNet and 3RScan with extra 3D QA data from two sources: (1) LEO [34], which provides 94k QA samples from 3RScan; and (2) MMScan [54], which offers 575k QA samples from ScanNet and 3RScan. The resulting checkpoint is denoted as “+QA”. Notably, these two datasets exhibit a highly repetitive and simplistic answer distribution. To quantify this, we normalize the answers by replacing common entities with placeholders (e.g., [COLOR]), transforming them into answer templates. As shown in Section 4.3, 39.2% of answers in “+QA” fall into the top-15 frequent templates, compared to only 12.8% in our curated data, demonstrating the simplistic distribution of “+QA”. The results in Section 4.3 reveal that despite incorporating substantially more QA data, “+QA” shows degradation in the overall performance compared to “Scan+3R”. This confirms that scaling with low-quality data can be detrimental, reinforcing our core curation principle: prioritizing data quality over sheer scale is essential for effective training of 3D-VL models.
@@
To probe the scaling effects, we ablate the proportions of training data based on full-domain coverage, specifically training with 12.5%, 25%, 50%, and 100% of the dataset. We report the overall performance by averaging metrics across all evaluated benchmarks. As shown in Fig. 7, the results exhibit a steady upward trend, demonstrating consistent improvements as the data scale increases. This highlights a desirable scaling behavior enabled by our data design, echoing the noisy trend or even degradation in performance when scaling with low-quality 3D-VL data. The consistent scaling effects underscore the importance of data curation in scaling for 3D-VL generalist training. Furthermore, the continued upward trend with no signs of saturation suggests that further performance gains may be achievable by scaling up more curated 3D-VL data.
@@ 4.4 Post-training
We explore the effect of post-training objective based on LEO-VL after instruction tuning. Specifically, we continue to fine-tune LEO-VL on SQA3D with SFT and SceneDPO, respectively. To construct negative answers , we first let LEO-VL generate predictions on the training set of SQA3D, and then employ GPT-4o [60] to rephrase correct predictions into hard negative samples. These rephrased negatives, along with the original incorrect predictions, form the negative answer set for SceneDPO. For negative scenes , we randomly sample a different scene from the training set. We evaluate the performance by two metrics: in-domain performance, computed as the average of EM and EM-R on SQA3D; and overall performance, averaged across all benchmarks. We remove the reference model in our implementation, given the empirical observation that it has a negligible impact while largely reducing resource costs. We set , and train both SFT and SceneDPO for 10 epochs.
Data | SQA3D | Avg. |
LEO-VL | 61.2 | 55.0 |
SFT | 62.5 | 53.7 |
SceneDPO | 62.7 | 53.8 |
Results.
The results are reported in Section 4.4 and Fig. 8. As shown in Section 4.4, SceneDPO improves both the in-domain performance on SQA3D and the overall performance across benchmarks, demonstrating its advantage in enhancing post-training performance. Fig. 8 shows consistent improvements in the accuracy over both answer and scene, indicating the acquisition of discriminative ability for the model. Notably, the significant gain in the accuracy over scene suggests the model’s evolution in addressing the issue of visual ignorance. On the other hand, the initially low accuracy also reveals the model’s limited ability in distinguishing semantically similar answers and correctly leveraging scene information for QA. This observation reinforces the need for effective post-training strategies like SceneDPO in enhancing the robustness of 3D-VL models.

### v2

@@ 3.3 Post-training
We argue that only SFT can be insufficient to build robust 3D VLMs, as evidenced by prior studies (Li et al., 2024b; Pi et al., 2024; Wang et al., 2024a). In particular, given the pronounced risk of overfitting in 3D VLMs (Deng et al., 2024; Huang et al., 2025b), it is crucial to design an effective post-training objective to enhance their robustness.
To this end, we propose SceneDPO, a novel post-training objective for 3D VLMs. We start with a DPO-like loss term that contrasts between positive answer and negative answer . Motivated by the issue of visual ignorance (Huang et al., 2025b), we introduce a loss term that contrasts between positive scene and negative scene . This discourages the model from predicting the current answer when conditioned on irrelevant scenes. To mitigate degradation of positive answers, we incorporate a negative log-likelihood loss term , which proves critical in prior works (Liu et al., 2024; Pang et al., 2024; Wang et al., 2024a). Given a training tuple , the overall loss is defined as follows:
@@ 4 Experiment
We first compare LEO-VL against state-of-the-art models on various 3D QA benchmarks in Section˜4.1, highlighting its performance and efficiency. In Section˜4.2, we present ablation studies on model design to show the effectiveness of CFG. In Section˜4.3, we analyze various training data configurations to reveal the influence of task diversity, scene domain coverage, data quality, and data scale on 3D-VL learning. In Section˜4.4, we further explore the post-training for 3D VLM.
@@
We initialize the 2D visual encoder and LLM with Qwen2.5-VL-7B-Instruct (Bai et al., 2025). The learnable parameters include position embedding, anchor embedding, and LoRA parameters (Hu et al., 2022) of the LLM, which amount to 66M in total. We set the voxel size to 0.2 meters and retain up to 750 scene tokens for CFG. We train LEO-VL on our comprehensive 3D-VL dataset for 5 epochs, which takes 2 days with 8 NVIDIA A100 80G GPUs. We adopt AdamW optimizer (Loshchilov & Hutter, 2017) with a base learning rate at , scheduled with linear warmup and cosine decay. We provide detailed hyperparameters in Appendix.
@@
Our evaluation mainly involves 3D QA, including ScanQA (Azuma et al., 2022) for general QA, SQA3D (Ma et al., 2023) for situated QA, MSQA (Linghu et al., 2024) for situated QA across multiple scene domains, and Beacon3D (Huang et al., 2025b) for robust QA evaluation in diverse scene domains. For the captioning task, we report the evaluation on Scan2Cap (Chen et al., 2021) in Appendix, as we think existing captioning benchmarks fall short in reflecting natural descriptive capabilities. Evaluation of other tasks is not reported due to the limited availability of standardized benchmarks. We report n-gram metrics (CIDEr and BLEU-4) for ScanQA, exact-match accuracy (EM and EM-R) for SQA3D, and GPT-Score for MSQA and Beacon3D.
@@ 4.1 Comparison with State-of-the-art Models
For comparison, we include state-of-the-art 3D VLMs across four categories: (1) query-based methods, ranging from early work like ScanQA (Azuma et al., 2022) to recent 3D-LLaVA (Deng et al., 2025); (2) object-centric methods, from 3D-VisTA (Zhu et al., 2023) to Inst3D-LMM (Yu et al., 2025); (3) voxel-based methods, such as Scene-LLM (Fu et al., 2025) and LLaVA-3D (Zhu et al., 2024a); and (4) video-based methods, including Video-3D LLM (Zheng et al., 2025) and GPT4Scene (Qi et al., 2025). We also show their scene representations and associated costs. For Scene-LLM, only the voxel size (0.18 meters) is available, which we estimate to yield a similar number of tokens as LLaVA-3D (voxel size 0.2 meters). For video-based methods, we list the number of frames, noting that 32 frames typically amount to over 6k scene tokens.
@@ Results.
As shown in Sections˜4.1 and 4.2, LEO-VL achieves state-of-the-art performance on most 3D QA benchmarks with a much lower representation cost. While video-based models show competitive performance, their representations involve thousands of tokens and limit data scalability, hindering scalable 3D-VL learning across more scene domains. In contrast, the efficient representation of LEO-VL enables 3D-VL data scaling for consistently strong performance across domains, including ScanNet, 3RScan, and MultiScan. These results underscore the advantages of our representation in efficiency, perception capability, and data scalability.
| Model | Scene (#tokens) | ScanQA (val) | SQA3D (test) | Beacon3D | Beacon3D | Beacon3D | |||||
| C | B-4 | EM | EM-R | Case | Obj. | Case | Obj. | Case | Obj. | ||
| ScanQA (Azuma et al., 2022) | Query (256) | 64.9 | 10.1 | 47.2 | - | - | - | - | - | - | - |
| 3D-LLM (Hong et al., 2023) | Query (32) | 74.5 | 12.9 | 49.8 | - | - | - | - | - | - | - |
| PQ3D (Zhu et al., 2024b) | Query (80) | - | - | 47.1 | - | 35.9 | 4.2 | 25.7 | 0.7 | 20.8 | 0.6 |
| DSPNet (Luo et al., 2025) | Query (256) | - | - | 50.4 | - | - | - | - | - | - | - |
| 3D-LLaVA (Deng et al., 2025) | Query (100) | 92.6 | 17.1 | 54.5 | 56.6 | - | - | - | - | - | - |
| 3D-VisTA (Zhu et al., 2023) | Object (80) | 69.6 | 10.4 | 48.5 | - | 43.2 | 7.3 | 25.7 | 3.3 | 19.1 | 0.0 |
| LEO (Huang et al., 2024b) | Object (60) | 101.4 | 13.2 | 50.0 | 52.4 | 45.2 | 7.5 | 44.0 | 1.5 | 26.2 | 0.6 |
| SceneVerse (Jia et al., 2024) | Object (80) | - | - | 49.9 | - | 40.5 | 4.7 | 37.4 | 0.4 | 28.9 | 3.1 |
| Chat-Scene (Huang et al., 2024a) | Object (200) | 87.7 | 14.3 | 54.6 | 57.5 | 49.8 | 10.9 | - | - | - | - |
| Inst3D-LMM (Yu et al., 2025) | Object (200) | 88.6 | 14.9 | - | - | - | - | - | - | - | - |
| Scene-LLM (Fu et al., 2025) | Voxel (0.18m) | 80.0 | 11.7 | 53.6 | - | - | - | - | - | - | - |
| LLaVA-3D (Zhu et al., 2024a) | Voxel (3096) | 91.7 | 14.5 | 55.6 | 57.6 | 59.1 | 19.0 | - | - | - | - |
| Video-3D LLM (Zheng et al., 2025) | Video (32 frames) | 100.5 | 16.3 | 57.7 | - | 59.0 | 17.9 | - | - | - | - |
| GPT4Scene (Qi et al., 2025) | Video (32 frames) | 96.3 | 15.5 | 59.4 | 62.4 | 57.2 | 17.9 | - | - | - | - |
| LEO-VL | Grid (750) | 100.4 | 15.5 | 60.8 | 63.7 | 59.5 | 19.2 | 48.2 | 3.3 | 37.7 | 6.9 |
@@ 4.2 Model Analysis
| Model | Count. | Exist. | Attr. | Spatial | Navi. | Others | Overall |
| GPT-4o∗‡ (OpenAI, 2023) | 32.3 | 79.3 | 79.0 | 37.0 | 31.7 | 91.6 | 52.3 |
| LEO (Huang et al., 2024b) | 32.5 | 88.5 | 58.7 | 44.2 | 39.6 | 81.4 | 54.8 |
| MSR3D (Linghu et al., 2024) | 32.3 | 93.1 | 50.0 | 46.5 | 54.1 | 75.6 | 54.2 |
| SplatTalk†‡ (Thai et al., 2025) | 19.6 | 60.3 | 44.0 | 35.8 | 35.5 | 61.8 | 41.8 |
| LEO-VL (ScanNet) | 32.0 | 92.7 | 59.2 | 57.3 | 48.1 | 83.5 | 58.4 |
| LEO-VL (All) | 39.3 | 92.7 | 56.9 | 59.3 | 59.7 | 82.8 | 61.7 |
@@
In Fig.˜4, we present statistics of the number of voxel tokens and CFG tokens on ScanNet. In addition to the distribution of token counts, we report two metrics: compression rate, defined as the ratio of CFG tokens to raw voxel tokens before condensation; and preservation rate, defined as the proportion of voxels retained by the CFG tokens. The results show that CFG achieves an average compression rate of about and an average preservation rate close to 99%. This demonstrates the high efficiency of our CFG representation.
@@
We provide a joint visualization of efficiency and accuracy in Fig.˜3. Accuracy is measured using the EM on SQA3D, the metric with the most available reference results. As shown in Fig.˜3, LEO-VL achieves a Pareto optimum between efficiency and accuracy. To investigate the efficiency of CFG, we ablate the representation with two alternatives on the ScanNet subset: voxel-full, which retains all voxel tokens without condensation; and voxel-sample, which downsamples voxel tokens to match the token count of CFG. The results in Fig.˜3 show that compared to voxel-full, voxel-sample suffers a drop in accuracy, while CFG even slightly improves the accuracy. These results suggest that high accuracy in 3D-VL tasks can be achieved with significantly lower token costs, as enabled by our efficient CFG representation.
@@
A key design of CFG lies in the disentanglement of 3D spatial modeling across vertical and horizontal directions. We train LEO-VL on the ScanNet subset under two ablated settings: “w/o z-pos”, where the z-axis position embedding is removed; and “w/o xyz-pos”, where both z-axis and xy-plane position embeddings are removed. As shown in Fig.˜6, both types of position embeddings play critical roles in 3D-VL tasks. Removing the z-axis position embedding leads to a general performance drop, while removing the xy-plane position embedding especially degrades performance on SQA3D. These results validate the effectiveness of our design in 3D spatial modeling.
@@ 4.3 Data Analysis
We conduct ablative studies on data to answer the following questions: (1) What is the effect of captioning, planning and dialogue tasks in 3D-VL learning? (2) What is the benefit of scaling across diverse scene domains? (3) What happens when we naively scale up 3D-VL data without considering data quality? (4) How effective is data scaling for LEO-VL with curated data?
@@
Based on LEO-VL (ScanNet), we ablate the training data with different task configurations: “w/o cap” removes both object captioning and scene captioning tasks, while “w/o pl+dl” removes planning and dialogue tasks. As shown in Fig.˜7, both task categories contribute meaningfully to the capabilities of LEO-VL. Captioning proves particularly beneficial for SQA3D, whereas planning and dialogue tasks are more influential for ScanQA. We further qualitatively test the “w/o cap” model to perform the scene captioning task. The model struggles to generate meaningful scene descriptions, likely due to the lack of training in captioning. These findings underscore the importance of task diversity for both acquiring task-specific skills and enhancing cross-task performance.
@@
To assess the impact of scaling across scene domains, we add a checkpoint trained on ScanNet and 3RScan (“Scan+3R”). As shown in Section˜4.3, “Scan+3R” yields significant improvements on 3RScan and ARKitScenes compared to “ScanNet”. Furthermore, “All” shows consistent improvements on all domains compared to “Scan+3R”. These results suggest that the performance gains stem from the training in more diverse domains, emphasizing the importance of cross-domain scaling for 3D-VL learning.
@@
Based on the results of “Scan+3R”, we explore the outcome of naively scaling up 3D-VL data without considering data quality. We expand our default subsets on ScanNet and 3RScan with extra 3D QA data (denoted as “QA”) from two sources: (1) LEO (Huang et al., 2024b), which provides 94k QA samples from 3RScan; and (2) MMScan (Lyu et al., 2024), which offers 575k QA samples from ScanNet and 3RScan. The resulting checkpoint is denoted as “+QA”. We present QA data statistics, especially the occupancy of top-15 frequent answer templates in Section˜4.3, which indicates a more simplistic answer distribution in “QA”. The results in Section˜4.3 reveal that despite substantially more QA data, “+QA” shows degradation in the overall performance compared to “Scan+3R”. This shows the harm of scaling with low-quality data, and suggests that prioritizing data quality over scale is critical for effective scaling of 3D-VL learning.
@@
We demonstrate that our curated data scheme exhibits consistent scaling effects. Specifically, we take 12.5%, 25%, and 50% of our full data for training, and report the overall performance by averaging metrics across all benchmarks. As shown in Fig.˜5, the results exhibit a steady upward trend, demonstrating consistent gains as the data scale increases. This reflects the desired scaling behavior enabled by our curated data scheme, in contrast to the degradation when scaling with low-quality data. Notably, the upward trend shows no sign of saturation, suggesting the potential of further scaling with curated 3D-VL data.
| Data | SQA3D | Beacon3D | Beacon3D |
| Baseline | 61.2 | 26.5 | 21.9 |
| SFT | 62.5 | 27.5 | 22.3 |
| GRPO | 51.6 | 28.1 | 25.0 |
| SceneDPO | 62.7 | 28.4 | 25.2 |
@@ 4.4 Post-training
We explore post-training for 3D VLM by continuing to train LEO-VL on SQA3D, comparing SceneDPO with SFT and GRPO. The evaluation includes in-domain test on SQA3D (ScanNet), and out-of-domain test on Beacon3D (3RScan and MultiScan). We report the average of EM and EM-R for SQA3D, and the average of case-centric and object-centric scores for Beacon3D. Training hyperparameters follow the instruction tuning stage. Specific implementations are as follows:
SFT. The implementation is the same as the previous instruction tuning stage.
GRPO. We design an accuracy reward that yields for EM correct, for EM-R correct, and otherwise. We have not included format reward due to the lack of cold start Chain-of-Thought (CoT) data. The group size is set to , clip range , and KL coefficient .
SceneDPO. We compile negative answers by bootstrapping on model’s predictions. We keep the incorrect ones and employ GPT-4o to rephrase the correct ones into negative samples. During post-training, we randomly sample a different scene as the negative scene . We set , , , and .
Quantitative Results.
As shown in Section˜4.3, all strategies outperform the baseline except GRPO’s performance on SQA3D. This implies that rewards may be too sparse to provide effective learning signals for 3D VLM on complex spatial reasoning tasks (e.g., SQA3D). However, GRPO exhibits better out-of-domain performance than SFT, indicating its comparative advantage in enhancing generalizability. In contrast, SceneDPO achieves the best for both in-domain and out-of-domain performances, highlighting its strength as a comprehensive objective. These findings suggest that the potential of GRPO-like post-training may be bottlenecked by the lack of 3D-VL CoT reasoning data, and SceneDPO can be a more practical objective for the post-training of current 3D VLMs.
Qualitative Results.
Fig.˜8 visualizes the dynamic curves of contrast accuracy across answers and scenes during post-training. We observe consistent increasing trends for both accuracies, reflecting an enhancement in the model’s ability to discriminate between positive and negative pairs. Notably, the initially low accuracy over scenes reveals the model’s limited ability in exploiting scene context for QA, i.e., the issue of visual ignorance. These observations resonate with our design motivation and demonstrate the efficacy of SceneDPO for enhancing the robustness of 3D VLMs.